OpenStudy (anonymous):

If y=2/3x, which of the following is an expression for {{y{{ in terms of x? For any x, {{X{{=2/3x squared

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x=36

OpenStudy (anonymous):

2/3 times=24

OpenStudy (anonymous):

so

OpenStudy (anonymous):

y=24

OpenStudy (anonymous):

huh..

OpenStudy (anonymous):

so, is {{y{{ 24?

hartnn (hartnn):

question unclear :(

OpenStudy (anonymous):

3 y = 2 x?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

for any number x, #x=2/3x squared

OpenStudy (anonymous):

y=2/3x

OpenStudy (anonymous):

Which of the following is an expression for #y in terms of x?

OpenStudy (anonymous):

2/3x cubed

OpenStudy (anonymous):

4/9 x squared

OpenStudy (anonymous):

4/9x cubed

OpenStudy (anonymous):

8/27 x squared

OpenStudy (anonymous):

8/ 27 cubed

OpenStudy (anonymous):

8/27 x cubed

OpenStudy (anonymous):

What is the #x

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

good question

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

let us say that x=36

OpenStudy (anonymous):

plug in 36 for 2/3x squared

OpenStudy (anonymous):

2/3 times 1,296

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

#x=864

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so

OpenStudy (anonymous):

let us do the same for y

OpenStudy (anonymous):

2/3 times 36=24

OpenStudy (anonymous):

#y=24

OpenStudy (anonymous):

#x=864

OpenStudy (anonymous):

ok so that means #y = 2/3y^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so, which of the following is an expression for #y in terms of x?

OpenStudy (anonymous):

then you substitute in y

OpenStudy (anonymous):

which is 2/3x

OpenStudy (anonymous):

2/3*(2/3x)^2

OpenStudy (anonymous):

8/27x^2

OpenStudy (anonymous):

well, they want you to find 24 in terms of x

OpenStudy (anonymous):

x=36

OpenStudy (anonymous):

so, essentially, which equation gives us 24, #y

OpenStudy (anonymous):

I thought i answered the question

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Somebody verify for me because i am getting confuse.. lol

OpenStudy (anonymous):

ill come back

OpenStudy (anonymous):

well

OpenStudy (anonymous):

the answer is D.

OpenStudy (anonymous):

but

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

D equals 384, not 24

OpenStudy (anonymous):

unless

OpenStudy (anonymous):

you plug in #y for x

OpenStudy (anonymous):

sp

OpenStudy (anonymous):

so

OpenStudy (anonymous):

2/3 times 24 squared

OpenStudy (anonymous):

2/3 times 576

OpenStudy (anonymous):

384

OpenStudy (anonymous):

8/27 x squared=384

OpenStudy (anonymous):

so you plug in #y for x in the first equation, and you then see which equation listed equals your result

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

next problem

OpenStudy (anonymous):

Lol wasn't that what i gave you

OpenStudy (anonymous):

xD